COMPARISON OF SPECTROSCOPIC AND AB INITIO STRUCTURES FOR THE HYDROGEN-BONDED COMPLEX TRIMETHYLAMINE-HYDROGEN SULFIDE

\underline{MICHAEL J. TUBERGEN}, and JONATHAN E. FLAD, Department of Chemistry, Kent State University, Kent, OH 44242; JANET E. DEL BENE, Department of Chemistry, Youngstown State University, Youngstown, OH 44555.

Rotational spectra have been recorded for six isotopomers of the trimethylamine-hydrogen sulfide complex using a Fouriertransform microwave spectrometer. The spectra were found to be characteristic of a symmetric top, (B + C)/2 = 1395.463 (1) MHz, and are indicative of free internal rotation of trimethylamine within the complex. A structure with a single, linear hydrogen bond (R_{N−S} = 3.36 (5) Åbest reproduces the moments of inertia of the six isotopic species, including three distinct deuterated complexes. The experimental structure is compared to the ab initio structure optimized at the MP2/6-31G(d,p) level, which predicts R_{N−S} = 3.328 Å. MP2/aug′-cc-pVTZ calculations were used to determine the binding energy of the complex and the barrier to an internal tunneling motion which exchanges the two H_{2}S hydrogens.